OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-2, 0) and (2, 8)? give a medal to first that answers

OpenStudy (anonymous):

Ok Lets see here

OpenStudy (anonymous):

First of all, you need to know the point-slope form equation. point-slope form: y1 - y2 = m (x1 - x2) Then, find the slope of the line. slope equation: y1 - y2 / x1 - x2 8 - 3 / 2 - 3 5 / -1 -5 Next, plug in the coordinates in the right place. x2: 2 or 3 y2: 8 or 4 plug-in: y - 4 = -5 (x - 3) y - 8 = -5 (x - 2) And that's the equation you can use either coordinates if you'd like.

OpenStudy (anonymous):

is that good enough lol

OpenStudy (anonymous):

what do you think the answer is

OpenStudy (anonymous):

i know what the answer is

OpenStudy (anonymous):

jsut wandering what you think it is

OpenStudy (anonymous):

ANSWER: y - 8 = -4(x - 2) or y - 4 = -4(x - 3)

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

wheres my medal

OpenStudy (anonymous):

these are the possible answers a. y = 2(x + 2) b. y = 2(x - 2) c. y = -2(x - 2) d. y = -2(x + 2)

OpenStudy (anonymous):

thats not the answer

OpenStudy (anonymous):

ooooo ok got you one moment

OpenStudy (anonymous):

Start by using the slope intercept form y1-y2=m(x1-x2) Then, leave y1 and x1 the same and plug in one of the coordinates: (-2,0) y-0=m(x+2) Next find m. m=slope. slope= change of y over change of x 0-8/-2-2 = -8/-4 = 2 your slope is 2. Plug that in for m. y-0=2(x+2) Simplify so you are in point slope form. y=2(x+2) A is your answer.

OpenStudy (anonymous):

ya nice job, a medal and a fan for you, fan me back